# Stong order $1$ adaptive approximation of jump-diffusion SDEs with discontinuous drift

### v1

@@ 3.2 Occupation time estimates
The proof of the following two Lemmas combines ideas of [57, Lemma 3.9] and [65, Lemma 6]. In a first step, we estimate the local time of $Z^{(\delta)}$. For this we denote for all $a\in{\mathbb{R}}$ by $(L_{t}^{a}(Z^{(\delta)}))_{t\in[0,T]}$ the local time of $Z^{(\delta)}$ in $a$.
@@
Let Assumption 3.1 hold. Then there exists a constant $c_{4}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$, and all $a\in{\mathbb{R}}$,
@@
By [63, Lemma 158] for all $a\in{\mathbb{R}}$,
@@
Here we denote $\operatorname{sgn}(x)=\mathds{1}_{(0,\infty)}(x)-\mathds{1}_{(-\infty,0]}(x)$ for all $x\in{\mathbb{R}}$. By (3.28) we obtain
@@
$\leq\Big{|}\big{|}Z^{(\delta)}_{t}-a\big{|}-\big{|}\widetilde{\xi}-a\big{|}\Big{|}+\Big{|}\int_{0}^{t}\operatorname{sgn}\big{(}Z^{(\delta)}_{u-}-a\big{)}\mathop{}\!\mathrm{d}Z^{(\delta)}_{u}\Big{|}$
@@
Estimating the expectation of each summand on the right hand side of (3.29) separately, we obtain for the first one using Lemma 3.9 (3.24) that there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that
@@
For the second summand of (3.29) we obtain by (3.11)
$\Big{|}\int_{0}^{t}\operatorname{sgn}\big{(}Z^{(\delta)}_{u-}-a\big{)}\mathop{}\!\mathrm{d}Z^{(\delta)}_{u}\Big{|}$ (3.31)
@@
Using Cauchy-Schwarz inequality, Itô isometry, Lemma 3.7 (3.22), and Lemma 3.9 (3.24) we obtain
@@
Using Cauchy-Schwarz inequality, Lemma 3.6, and Lemma 3.9 (3.24) we obtain that there exists $\hat{c}\in(0,\infty)$ such that
@@
By (3.31), (3.32), (3.33), and (3.34) there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that
@@
For the expectation of the third summand of (3.29) we apply Cauchy-Schwarz inequality, Lemma 3.6, and Lemma 3.9 (3.24) to obtain
@@
By (3.29), (3.30), (3.35), and (3.36) we obtain that there exists a constant $c_{4}\in(0,\infty)$ such that
@@
Let Assumption 3.1 hold, let $f\colon[0,\infty)\to[0,\infty)$ be a ([0,))/$\mathcal{B}([0,\infty))/\mathcal{B}([0,\infty))$-measurable and bounded function, and let $\gamma\in(0,3/2)$. Then there exists a constant $c_{5}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $\varepsilon\in(0,\varepsilon_{0}]$,
@@
Hence we fix $k\in\{1,\ldots,\widetilde{m}\}$ and assume $\delta\in(0,\delta_{0}]$. Since $\widetilde{\sigma}(\zeta_{k})\neq 0$, there exist $\widetilde{\varepsilon},\widetilde{c}_{2}\in(0,\infty)$ such that
@@
and we obtain for all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon_{0}]$
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\Big{(}\widetilde{\sigma}^{2}\big{(}Z^{(\delta)}_{t}\big{)}-\Big{(}\sum_{n=0}^{\infty}\Big{(}\widetilde{\sigma}\big{(}Z^{(\delta)}_{\tau_{n}}\big{)}+\int_{\tau_{n}}^{t}\widetilde{\sigma}\big{(}Z^{(\delta)}_{\tau_{n}}\big{)}d_{\widetilde{\sigma}}\big{(}Z^{(\delta)}_{\tau_{n}}\big{)}\mathop{}\!\mathrm{d}W_{s}\Big{)}\mathds{1}_{(\tau_{n},\tau_{n+1}]}(t)\Big{)}^{2}\Big{)}\mathop{}\!\mathrm{d}t\Big{]}$
@@
see [63, Theorem 88]. Hence we use that $x\mapsto f(|x-\zeta_{k}|)\mathds{1}_{(\zeta_{k}-\varepsilon,\zeta_{k}+\varepsilon)}(x)$ is a Borel-measurable and bounded function, apply [63, Lemma 159] and Lemma 3.27 to obtain
@@
Using $|a^{2}-b^{2}|=|a-b|\cdot|a+b|$ for all $a,b\in{\mathbb{R}}$ yields
@@
and apply Lemma 3.7, and Lemma 3.9 (3.24) and (3.25) to obtain
@@
which together with Lemma 3.7, and Lemma 3.9 (3.24) yields that there exists a constant $\widetilde{c}_{3}\in(0,\infty)$
@@
Plugging (3.47) and (3.49) into (3.45) yields that there exists a constant $\widetilde{c}_{4}\in(0,\infty)$ such that
@@
Next we use (3.44), Hölder’s inequality with $q\in(1,\infty)$, and (3.50) in (LABEL:pOTE27) to obtain for all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon_{0}]$,
@@
Iterating this and then estimating ${\mathbb{P}}(|Z^{(\delta)}_{t}-\zeta_{k}|\leq\varepsilon)\leq 1$ we obtain
@@
Plugging (3.54) in (3.51) we get that there exists a constant $\widetilde{c}_{7}\in(0,\infty)$ such that
@@
By Youngs in equality we have
@@
Hence we obtain that there exists a constant $\widetilde{c}_{8}\in(0,\infty)$ such that
@@
This implies the claim for all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon_{0}]$ as $\frac{3}{2q}\uparrow\frac{3}{2}$ for $q\downarrow 1$ and $\frac{1}{2}+\frac{1}{2q}+\frac{1}{2q^{2}}\uparrow\frac{3}{2}$ for $q\downarrow 1$. This is enough since in case that $\widetilde{\varepsilon}<\varepsilon_{0}$ for $\varepsilon\in(\widetilde{\varepsilon},\varepsilon_{0}]$ we calculate
@@
Here $C([0,\infty),{\mathbb{R}})$ is the space of all continuous functions from $[0,\infty)$ to ${\mathbb{R}}$. From [65, Proof of Lemma 7] we know the following three estimates. There exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $y\in\Theta^{\varepsilon_{2}^{\delta}}$,
@@
Let Assumption 3.1 hold. Then for all $p\in{\mathbb{N}}_{0}$, $q\in{\mathbb{N}}$, $u\in[0,T]$, and $\delta\in(0,\delta_{0}]$ there exists a constant $c_{7}\in(0,\infty)$ such that
@@
Since ${\mathbb{R}}=(\Theta^{\varepsilon_{0}})^{c}\cup(\Theta^{\varepsilon_{0}}\setminus\Theta^{\varepsilon_{1}^{\delta}})\cup(\Theta^{\varepsilon_{1}^{\delta}}\setminus\Theta^{\varepsilon_{2}^{\delta}})\cup\Theta^{\varepsilon_{2}^{\delta}}$, we split
@@
For the first summand we note that
@@
Using this, Hölder’s inequality, Markov’s inequality, and Lemma 3.9 we obtain
@@
For the second summand we calculate using Lemma 3.9, Lemma 3.12 (iii) with $\alpha=1$ and $q=2p$ that
@@
Similar for the third summand we calculate using Lemma 3.9, Lemma 3.12 (ii) with $\alpha=1$ and $q=2p$ that
@@
For the fourth summand we calculate using Hölder’s inequality
@@
Further we calculate
@@
For all $n\in{\mathbb{N}}$, and all $\omega\in\{Z_{\underline{t}}^{(\delta)}\in\Theta^{\varepsilon_{2}^{\delta}}\}$ we have $t-\underline{t}\leq\delta^{2}\log^{4}(\delta^{-1})$. Using this and Lemma 3.9 we calculate
${\mathbb{E}}\Big{[}\big{|}\widetilde{\mu}(Z^{(\delta)}_{\underline{t}})(t-\underline{t})\big{|}^{4p}\mathds{1}_{\Theta^{\varepsilon_{2}^{\delta}}}(Z_{\tau_{n}}^{(\delta)})\Big{]}\leq{\mathbb{E}}\Big{[}c_{\widetilde{\mu}}^{4p}2^{4p-1}(1+|Z^{(\delta)}_{\underline{t}}|^{4p})\delta^{8p}\log^{16p}(\delta^{-1})\mathds{1}_{\Theta^{\varepsilon_{2}^{\delta}}}(Z_{\underline{t}}^{(\delta)})\Big{]}$ (3.76)
@@
For the second summand we use Lemma 3.13 to obtain
@@
For the third summand we combine the previous arguments to obtain
@@
Hence combining (3.75), (3.76), (3.77) and (3.78) there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that
@@
Further we calculate using Jensen’s inequality, Lemma 3.11 with $f\equiv 1$ and $\gamma=1/2$ and Lemma 3.12 (i) with $\alpha=1$ and $q=1$ that there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that
@@
Plugging (3.79) and (3.80) in (3.74) we get that there exists $\widetilde{c}_{3},\widetilde{c}_{4}\in(0,\infty)$

### v2

@@ 3.2 Occupation time estimates
To prove occupation time estimates, we combine and adapt strategies presented in [57] and [65]. While the results and proofs of Lemma 3.11, Lemma 3.12, and Lemma 3.14 are closely based on the corresponding lemmas in [65], Lemma 3.10 and Lemma 3.13 present results with proofs heavily influenced by the additional jump noise that make use of similar ideas as in [57].
In a first step, we estimate the local time of $Z^{(\delta)}$. For this we denote for all $a\in{\mathbb{R}}$ by $(L_{t}^{a}(Z^{(\delta)}))_{t\in[0,T]}$ the local time of $Z^{(\delta)}$ in $a$. The local time of SDE (3.1) is defined for all $t\in[0,T]$ and all $a\in{\mathbb{R}}$ by
$L_{t}^{a}(Z^{(\delta)})=2\Big($ $(Z_{t}^{(\delta)}-a)^{+}-(Z_{0}^{(\delta)}-a)^{+}-\int_{0}^{t}\mathds{1}_{\{Z^{(\delta)}_{u-}>a\}}\mathop{}\!\mathrm{d}Z^{(\delta)}_{u}$ (3.27)
$-\int_{0}^{t}\Big(\big(Z^{(\delta)}_{u-}+\widetilde{\rho}\big(Z^{(\delta)}_{u-}\big)-a\big)^{+}-\big(Z^{(\delta)}_{u-}-a\big)^{+}-\mathds{1}_{\{Z^{(\delta)}_{u-}>a\}}\widetilde{\rho}\big(Z^{(\delta)}_{u-}\big)\Big)\mathop{}\!\mathrm{d}N_{u}\Big),$
where $(\cdot)^{+}=\max\{\cdot,0\}$, see [63, Definition 156].
@@
Let Assumption 3.1 hold. Then there exists a constant $c_{4}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$, all $t\in[0,T]$ and all $a\in{\mathbb{R}}$,
@@
By [63, Lemma 158] for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
Here we denote $\operatorname{sgn}(x)=\mathds{1}_{(0,\infty)}(x)-\mathds{1}_{(-\infty,0]}(x)$ for all $x\in{\mathbb{R}}$. By (3.29) we obtain for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
[…]
@@
Estimating the expectation of each summand on the right hand side of (3.30) separately, we obtain for the first one using Lemma 3.9 (3.24) that there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
For the second summand of (3.30) we obtain by (3.11) for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
[…]
@@
Using Cauchy-Schwarz inequality, Itô isometry, Lemma 3.7 (3.22), and Lemma 3.9 (3.24) we obtain for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
Using Cauchy-Schwarz inequality, Lemma 3.6, and Lemma 3.9 (3.24) we obtain that there exists $\hat{c}\in(0,\infty)$ such that for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
By (3.32), (3.33), (3.34), and (3.35) there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
For the expectation of the third summand of (3.30) we apply Cauchy-Schwarz inequality, Lemma 3.6, and Lemma 3.9 (3.24) to obtain that for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
By (3.30), (3.31), (3.36), and (3.37) we obtain that there exists a constant $c_{4}\in(0,\infty)$ such that for all $a\in{\mathbb{R}}$, all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
Let Assumption 3.1 hold, let $f\colon[0,\infty)\to[0,\infty)$ be a $\mathcal{B}([0,\infty))/\mathcal{B}([0,\infty))$-measurable and bounded function, and let $\gamma\in(0,3/2)$. Then there exists a constant $c_{5}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $\varepsilon\in(0,\varepsilon_{0}]$,
@@
Hence we fix $k\in\{1,\ldots,\widetilde{m}\}$. Since $\widetilde{\sigma}(\zeta_{k})\neq 0$, there exist $\widetilde{\varepsilon},\widetilde{c}_{2}\in(0,\infty)$ such that
@@
and we obtain for all $\delta\in(0,\delta_{0}]$ and all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon_{0}]$,
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\Big(\widetilde{\sigma}^{2}\big(Z^{(\delta)}_{t}\big)-\Big(\sum_{n=0}^{\infty}\Big(\widetilde{\sigma}\big(Z^{(\delta)}_{\tau_{n}}\big)+\int_{\tau_{n}}^{t}\widetilde{\sigma}\big(Z^{(\delta)}_{\tau_{n}}\big)d_{\widetilde{\sigma}}\big(Z^{(\delta)}_{\tau_{n}}\big)\mathop{}\!\mathrm{d}W_{s}\Big)\mathds{1}_{(\tau_{n},\tau_{n+1}]}(t)\Big)^{2}\Big)\mathop{}\!\mathrm{d}t\Big].$
@@
see [63, Theorem 88]. Hence we use that $x\mapsto f(|x-\zeta_{k}|)\mathds{1}_{(\zeta_{k}-\varepsilon,\zeta_{k}+\varepsilon)}(x)$ is a Borel-measurable and bounded function, apply [63, Lemma 159] and Lemma 3.28 to obtain for all $\delta\in(0,\delta_{0}]$, all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon_{0}]$, and all $t\in[0,T]$,
@@
Using $|a^{2}-b^{2}|=|a-b|\cdot|a+b|$ for all $a,b\in{\mathbb{R}}$ yields for all $\delta\in(0,\delta_{0}]$ and all $u\in[0,t]$,
@@
and apply Lemma 3.7, and Lemma 3.9 (3.24) and (3.25) to obtain for all $\delta\in(0,\delta_{0}]$ and all $u\in[0,t]$,
@@
which together with Lemma 3.7, and Lemma 3.9 (3.24) yields that there exists a constant $\widetilde{c}_{3}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $u\in[0,t]$,
@@
Plugging (3.48) and (3.50) into (3.46) yields that there exists a constant $\widetilde{c}_{4}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $u\in[0,t]$,
@@
Next we use (3.45), Hölder’s inequality with $q\in(1,\infty)$, and (3.51) in (LABEL:pOTE27) to obtain for all $\delta\in(0,\delta_{0}]$ and all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon_{0}]$,
@@
Iterating this and then estimating ${\mathbb{P}}(|Z^{(\delta)}_{t}-\zeta_{k}|\leq\varepsilon)\leq 1$ we obtain for all $\delta\in(0,\delta_{0}]$ and all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon]$,
@@
Plugging (3.55) in (3.52) we get that there exists a constant $\widetilde{c}_{7}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon]$,
@@
By Young’s inequality we have for all $\delta\in(0,\delta_{0}]$ and all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon]$,
@@
Hence we obtain that there exists a constant $\widetilde{c}_{8}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon]$,
@@
This implies the claim for all $\varepsilon\in(0,\widetilde{\varepsilon}\wedge\varepsilon_{0}]$ as $\frac{3}{2q}\uparrow\frac{3}{2}$ for $q\downarrow 1$ and $\frac{1}{2}+\frac{1}{2q}+\frac{1}{2q^{2}}\uparrow\frac{3}{2}$ for $q\downarrow 1$. This is enough since in case that $\widetilde{\varepsilon}<\varepsilon_{0}$ for $\varepsilon\in(\widetilde{\varepsilon},\varepsilon_{0}]$ we calculate for all $\delta\in(0,\delta_{0}]$,
@@
Here $C([0,\infty),{\mathbb{R}})$ is the space of all continuous functions from $0,\infty)$ to $\mathbb{R}$}. From [65, Proof of Lemma 7] we know the following three estimates. There exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $y\in\Theta^{\varepsilon_{2}^{\delta}}$,
@@
Let Assumption 3.1 hold. Then for all $p\in{\mathbb{N}}_{0}$ and $q\in{\mathbb{N}}$ there exists a constant $c_{7}\in(0,\infty)$ such that for all $t\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
Since ${\mathbb{R}}=(\Theta^{\varepsilon_{0}})^{c}\cup(\Theta^{\varepsilon_{0}}\setminus\Theta^{\varepsilon_{1}^{\delta}})\cup(\Theta^{\varepsilon_{1}^{\delta}}\setminus\Theta^{\varepsilon_{2}^{\delta}})\cup\Theta^{\varepsilon_{2}^{\delta}}$, we split for all $\delta\in(0,\delta_{0}]$
@@
For the first summand we note that for all $t\in[0,T]$ and all $\delta\in(0,\delta_{0}]$,
@@
Using this, Hölder’s inequality, Markov’s inequality, and Lemma 3.9 we obtain for all $\delta\in(0,\delta_{0}]$,
@@
For the second summand we calculate using Lemma 3.9, Lemma 3.12 (iii) with $\alpha=1$ and $q=2p$ that for all $\delta\in(0,\delta_{0}]$,
@@
Similar for the third summand we calculate using Lemma 3.9, Lemma 3.12 (ii) with $\alpha=1$ and $q=2p$ that for all $\delta\in(0,\delta_{0}]$,
@@
For the fourth summand we calculate using Hölder’s inequality for all $\delta\in(0,\delta_{0}]$,
@@
Further we calculate for all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
For all $n\in{\mathbb{N}}$, and all $\omega\in\{Z_{\underline{t}}^{(\delta)}\in\Theta^{\varepsilon_{2}^{\delta}}\}$ we have $t-\underline{t}(\omega)\leq\delta^{2}\log^{4}(\delta^{-1})$. Using this and Lemma 3.9 we calculate for all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
${\mathbb{E}}\Big[\big|\widetilde{\mu}(Z^{(\delta)}_{\underline{t}})(t-\underline{t})\big|^{4p}\mathds{1}_{\Theta^{\varepsilon_{2}^{\delta}}}(Z_{\underline{t}}^{(\delta)})\Big]\leq{\mathbb{E}}\Big[c_{\widetilde{\mu}}^{4p}2^{4p-1}(1+|Z^{(\delta)}_{\underline{t}}|^{4p})\delta^{8p}\log^{16p}(\delta^{-1})\mathds{1}_{\Theta^{\varepsilon_{2}^{\delta}}}(Z_{\underline{t}}^{(\delta)})\Big]$ (3.77)
@@
For the second summand we use Lemma 3.13 to obtain for all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
For the third summand we combine the previous arguments to obtain for all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
Hence combining (3.76), (3.77), (3.78) and (3.79) there exists a constant $\widetilde{c}_{1}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$ and all $t\in[0,T]$,
@@
Further we calculate using Jensen’s inequality, Lemma 3.11 with $f\equiv 1$ and $\gamma=1/2$ and Lemma 3.12 (i) with $\alpha=1$ and $q=1$ that there exists a constant $\widetilde{c}_{2}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,
@@
Plugging (3.80) and (3.81) in (3.75) we get that there exists $\widetilde{c}_{3},\widetilde{c}_{4}\in(0,\infty)$ such that for all $\delta\in(0,\delta_{0}]$,